# Gravitational Wave with Domain Wall Dominance

### v1

@@ I Introduction
Topological defects become natural candidates. In fact, their existence and cosmological production are ubiquitous in theories where a global symmetry is spontaneously broken. In addition, the advent of generalized global symmetry [1] (see [2, 3, 4, 5, 6, 7] for recent reviews) and recent applications in particle physics [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] shed light on new theoretical perspectives manifested in the physics of topological defects, e.g., global structure of gauge groups [18, 17] or TQFT couplings [38, 13]. This also motivates us to explore the possible role of these defects in cosmology and universal/distinctive observational signatures associated with them in more detail.
@@ III.1 GW Emission Power
On the other hand, once the universe enters the DWD phase, the situation drastically changes. While there is only one relevant scale, $L\sim H^{-1}\sim t$, during the scaling regime in RD, there are three scales we need to consider in DWD: $L,H^{-1}$, and DW fluctuation scale, $H^{-2}/L$. Recalling that while $H^{-1}\sim t$, the characteristic length of DW scales as $L(t)\propto a(t)\sim t^{2}$ in DWD: $L$ increases faster than $H^{-1}$. Specifically, considering the regime where $L\gg H^{-1}(t)$ at $t>t_{\rm weq}$, the amplitude of the fluctuation is estimated to be $\sim H^{-2}/L$ (versus $\sim H^{-1}$ in the scaling regime), as schematically depicted in Fig. 2. This shows that DW fluctuations are suppressed by a factor $H^{-1}/L\ll 1$. Therefore, unlike the scaling regime, the quadrupole moment in a single Hubble patch is expected to be $Q\sim M_{w}(H^{-2}/L)^{2}\sim\sigma H^{-6}/L^{2}$, where the mass of the wall within the horizon is given to be $M_{w}\sim\sigma H^{-2}$. The total GW power emitted within a single Hubble volume is $P_{\rm GW}^{\rm(DWD)}\sim G\sigma^{2}H^{-6}/L^{4}$ assuming the oscillation frequency to be $\omega\sim H$. Therefore, we expect largely suppressed GW emission during DWD,
@@ III.2 Characteristic Scales and Transfer Functions
Figure 3: The frequencies $f_{\rm weq}=k_{\rm weq}/2\pi$ (black solid) and $f_{D}=k_{D}/2\pi$ (gray dashed) corresponding to the comoving horizon scale at $t_{\rm weq}$ and $t_{D}$, respectively. The gray-colored region corresponds to $t_{\rm D}<t_{\rm WD}$, in which DWD phase is not realized. The two colored (blue and red) dots denote two example benchmark points: (BM1) $T_{D}=10^{5}\,\text{GeV}$ and $\sigma=(10^{10}\,\text{GeV})^{3}$ and (BM2) $T_{D}=10^{11}\,\text{GeV}$ and $\sigma=(10^{15}\,\text{GeV})^{3}$. The green dot (BM3) $T_{D}=10^{5}\,\text{GeV}$ and $\sigma=(10^{9}\,\text{GeV})^{3}$ corresponds to a parameter set where the DWD phase does not occur. In the light gray region, the dotted lines denote the peak frequency determined by $T_{D}$. The green-colored region is ruled out due to the large inhomogeneities induced on the CMB/LSS. The blue- and orange-shaded regions are constrained by CMB [60] and PTA [61, 62], respectively.
@@ III.3.2 GW from DW with DWD
Spectra for $k<k_{r}$ show a strong dependence on the choice of $\alpha$, and these IR modes are also bounded by the causality constraint. The shape of the causality tail depends on the background evolution [79]. In deep IR regime, $k<k_{\rm weq}$, modes stay extensively on superhorizon and enter the horizon later during RD after DW decay. Therefore, these modes follow $k^{3}$ profile [80, 79]. On the other hand, modes with $k_{\rm weq}<k<k_{r}$ undergo a non-trivial evolution history, and the spectral shape also gets modified to $k^{-3}$, as explicitly shown in Sec. S-III.2. Fig. 4 shows that, for $\alpha>7/2$, the GW spectrum is dominated and lower-bounded by the causality tail from the GW emitted at $t=rt_{\rm weq}$.

### v2

@@ I Introduction
Topological defects become natural candidates. In fact, their existence and cosmological production are ubiquitous in theories where a global symmetry is spontaneously broken. In addition, the advent of generalized global symmetry [1] (see [2, 3, 4, 5, 6, 7] for recent reviews) and recent applications in particle physics [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] shed light on new theoretical perspectives manifested in the physics of topological defects, e.g., global structure of gauge groups [18, 17] or TQFT couplings [40, 13]. This also motivates us to explore the possible role of these defects in cosmology and universal/distinctive observational signatures associated with them in more detail.
@@ III.1 GW Emission Power
On the other hand, once the universe enters the DWD phase, the situation drastically changes. While there is only one relevant scale, $L\sim H^{-1}\sim t$, during the scaling regime in RD, there are three scales we need to consider in DWD: $L,H^{-1}$, and DW fluctuation scale, $H^{-2}/L$. Recalling that while $H^{-1}\sim t$, the characteristic length of DW scales as $L(t)\propto a(t)\propto t^{2}$ in DWD: $L$ increases faster than $H^{-1}$. Specifically, considering the regime where $L\gg H^{-1}(t)$ at $t>t_{\rm weq}$, the amplitude of the fluctuation is estimated to be $\sim H^{-2}/L$ (versus $\sim H^{-1}$ in the scaling regime), as schematically depicted in Fig. 2. This shows that DW fluctuations are suppressed by a factor $H^{-1}/L\ll 1$. Therefore, unlike the scaling regime, the quadrupole moment in a single Hubble patch is expected to be $Q\sim M_{w}(H^{-2}/L)^{2}\sim\sigma H^{-6}/L^{2}$, where the mass of the wall within the horizon is given to be $M_{w}\sim\sigma H^{-2}$. The total GW power emitted within a single Hubble volume is $P_{\rm GW}^{\rm(DWD)}\sim G\sigma^{2}H^{-6}/L^{4}$ assuming the oscillation frequency to be $\omega\sim H$. Therefore, we expect largely suppressed GW emission during DWD,
@@ III.2 Characteristic Scales and Transfer Functions
Figure 3: The frequencies $f_{\rm weq}=k_{\rm weq}/(2\pi)$ (black solid) and $f_{D}=k_{D}/(2\pi)$ (gray dashed) corresponding to the comoving horizon scale at $t_{\rm weq}$ and $t_{D}$, respectively. The gray-colored region corresponds to $t_{D}<t_{\rm weq}$, in which DWD phase is not realized. The two colored (blue and red) dots denote two example benchmark points: (BM1) $T_{D}=10^{5}\,\text{GeV}$ and $\sigma=(10^{10}\,\text{GeV})^{3}$ and (BM2) $T_{D}=10^{11}\,\text{GeV}$ and $\sigma=(10^{15}\,\text{GeV})^{3}$. The green dot (BM3) $T_{D}=10^{5}\,\text{GeV}$ and $\sigma=(10^{9}\,\text{GeV})^{3}$ corresponds to a parameter set where the DWD phase does not occur. In the light gray region, the dotted lines denote the peak frequency determined by $T_{D}$. The green-colored region is ruled out due to the large inhomogeneities induced on the CMB/LSS. The blue- and orange-shaded regions are constrained by CMB [62] and PTA [63, 64], respectively.
@@ III.3.2 GW from DW with DWD
Spectra for $k<k_{r}$ show a strong dependence on the choice of $\alpha$, and these IR modes are also bounded by the causality constraint. The shape of the causality tail depends on the background evolution [81]. In deep IR regime, $k<k_{\rm weq}$, modes stay extensively on superhorizon and enter the horizon later during RD after DW decay. Therefore, these modes follow $k^{3}$ profile [82, 83, 81]. On the other hand, modes with $k_{\rm weq}<k<k_{r}$ undergo a non-trivial evolution history, and the spectral shape also gets modified to $k^{-3}$, as explicitly shown in Sec. S-III.2. Fig. 4 shows that, for $\alpha>7/2$, the GW spectrum is dominated and lower-bounded by the causality tail from the GW emitted at $t=rt_{\rm weq}$.
@@ Acknowledgements.
The existence of DWs in the early universe can also generate scalar perturbations, potentially leading to large inhomogeneities, which may be constrained by the CMB or large-scale structure (LSS). In this section, we briefly discuss these aspects. (Fluctuations from DW networks could also result in primordial black hole production [89, 90, 91, 92, 93, 94, 95, 96]. It is of interest to understand how these considerations are modified in the DWD scenario.)